# Directions of Curvature as an Explanation for Loss of Plasticity

### v2

@@ 1 Introduction
A longstanding goal of machine learning research is to develop algorithms that can learn continually and cope with unforeseen changes in the data distribution (Ring, 1994; Thrun, 1998). Current learning algorithms, however, struggle to learn from dynamically changing targets and are unable to adapt gracefully to unforeseen changes in the distribution during the learning process (Zilly et al., 2021; Abbas et al., 2023; Lyle et al., 2023; Dohare et al., 2023a). Such limitations can be seen to be a byproduct of assuming, one way or another, the problem is stationary. Recently, there has been growing recognition of the fact that there are limitations to what can be learned from a fixed and unchanging dataset (Hoffmann et al., 2022) and that there are implicit non-stationarities in many problems of interest (Igl et al., 2021).
The concept of plasticity has been receiving growing attention in the continual learning literature, where the loss of plasticity—a reduction in the ability to learn new things (Dohare et al., 2023a; Lyle et al., 2023)—has been noted as a critical shortcoming in current learning algorithms. That is, learning algorithms that are performant in the non-continual learning setting, and more specifically neural networks, often struggle when applied to continual learning problems, exhibiting a striking loss of plasticity such that learning slows down or even halts after successive changes in the learning distribution. Such a loss of plasticity can be readily observed in settings where a neural network must continue to learn after changes occur in the observations or targets.
Several aspects of a learning algorithm have been found to contribute to, or mitigate, loss of plasticity. Examples include the optimizer (Dohare et al., 2023a), the step-size (Ash & Adams, 2020; Berariu et al., 2021), the number of updates (Lyle et al., 2023), the activation function (Abbas et al., 2023), and the use of specific regularizers (Dohare et al., 2021; Kumar et al., 2023; Lyle et al., 2021). Such factors hint that there might be simpler underlying mechanisms mechanisms for loss of plasticity. For example, the success of several methods that regularize neural networks towards properties of the initialization suggests that some property of the initialization mitigates loss of plasticity. Unfortunately, no such property has yet been identified. Some factors that have been found to correlate with loss of plasticity include, a decrease in the gradient or update norm (Abbas et al., 2023), neuron dormancy (Sokar et al., 2023), and an increase in the norm of the parameters (Nikishin et al., 2022).

### v4

@@ 1 Introduction
The goal of continual learning research is to develop algorithms that can continue to learn from a dynamic data distribution (Ring, 1994; Thrun, 1998). While current machine learning algorithms are capable of learning from a fixed dataset and generalizing to unseen data from the same distribution, these algorithms can struggle to adapt to changes in the distribution over the course of learning (Zilly et al., 2021; Abbas et al., 2023; Lyle et al., 2023; Dohare et al., 2023a). The ability to learn from new data, also referred to as plasticity, is one way in which algorithms can adapt to such changes. It has recently been noted that loss of plasticity—a reduced ability to learn new things (Dohare et al., 2023a; Lyle et al., 2023)—is a critical shortcoming of neural network learning algorithms in continual learning settings. In particular, neural network algorithms that are trained on a changing data distribution can lead to progressively slower learning, or completely stop learning altogether.
Identifying the mechanisms behind neural network plasticity is an active area of research, with several components potentially contributing to improving plasticity, or otherwise mitigating loss of plasticity. For example, the use of stateful optimizers have been found to exacerbate loss of plasticity due to inaccurate gradient estimates after a data distribution change (Dohare et al., 2023a; Lyle et al., 2023). Tuning other properties of the optimization process, such as the step-size (Ash & Adams, 2020; Berariu et al., 2021) and the number of updates (Lyle et al., 2023), have been found to mitigate loss of plasticity. Saturating activation functions can also lead to loss of plasticity by limiting capacity (Sokar et al., 2023), which can be mitigated by non-saturating activation functions (Abbas et al., 2023). While a full mechanistic understanding of plasticity has not yet been identified, these experimental results suggest that optimization dynamics play an important role in sustaining neural network properties needed for plasticity. Some properties that have been found to correlate with loss of plasticity include, a decrease in the gradient or update norm (Abbas et al., 2023), neuron dormancy (Sokar et al., 2023), and an increase in the norm of the parameters (Nikishin et al., 2022). Unfortunately, these properties do not explain loss of plasticity in all the situations that it occurs.